OpenStudy (anonymous):

need help how do i do this 1/2 / 1/3= im totally lost on this

OpenStudy (anonymous):

Brackets missing

OpenStudy (anonymous):

huh? no its dividing fractions

OpenStudy (anonymous):

i forgot the steps

OpenStudy (anonymous):

\[(1/2)/(1/3)=3/2\] and \[1/(2/(1/3))=1/6\]So, read question carefully.

OpenStudy (anonymous):

ok thanx